In 1990, out of a total of 2000 students in a college, 1400 were for graduation and the rest for post graduation (P.G). Out of 1400 graduates students, 100 were girls, however in all there were 600 girls in the college. In 1995, number of graduates students increase to 1700 out of which 250 were girls, but the number of P.G students fall on 500 of which only 50 were boys. In year 2000, out if 800girls, 650 were for graduation, whereas the total number of graduates was 2200. The number of boys and girls in P.G classes was equal. Represent the above information in a tabular form. Also calculate the percentage increase in the number of graduate students in 2000 as compared to 1990.
1
Expert's answer
2019-10-23T15:10:07-0400

"\begin{array}{c:c:c:c:c:}
 \begin{matrix}
 Year \\
 Degrees
\end{matrix} & 1990 & 1995 & 2000 & Total \\ \hline
 Graduation\ \ \ \ \ \ \ \ \ \ \ \ \ \ \begin{matrix}
 Boys \\
 Girls
\end{matrix} & \begin{matrix}
 1300 \\
 100
\end{matrix} & \begin{matrix}
 1450 \\
 250
\end{matrix}& \begin{matrix}
 1550 \\
 650
\end{matrix} & \begin{matrix}
 4300 \\
 1000
\end{matrix} \\
 \hdashline
 Total\ (a) & 1400 & 1700 & 2200 & 5300 
\\
 \hdashline
 Post-Graduation\ \begin{matrix}
 Boys \\
 Girls
\end{matrix} & \begin{matrix}
 100 \\
 500
\end{matrix} & \begin{matrix}
 50 \\
 450
\end{matrix}& \begin{matrix}
 150 \\
 150
\end{matrix} & \begin{matrix}
 300 \\
 1100
\end{matrix} 
\\
 \hdashline
 Total \ (b) & 600 & 500 & 300 & 1400 
\\
 \hdashline
 Total\ (a)+(b) & 2000 & 2200 & 2500 & 6700 
\end{array}"

Calculate the percentage increase in the number of graduate students in 2000 as compared to 1990.


"{(2200-1400) \over 1400}\times100\approx57.143\%"